Find the volume of a sphere with a 3inch radius(use 3.14)?
Find the volume of cone that is 9in high and has a 3inch radius?(use 3.14)
Find the area of a triangular piece of cloth that has a 10cm base and is 15cm high
Thank-U !! God Bless U

Answer by rfer(16322)

1)
V=Pid^3/6
V=3.14*6^3/6
V=113 cu cm
2)
V=Pir^2h/3
V=3.14*3^2*9/3
V=84.78 cu in
3)
A=l*w/2
A=10*15/2
A=75 sq cm